OpenStudy (logan13):

What is the solution to the equation –14 + x = 5? A. 9 B. 19 C. –9 D. 5

OpenStudy (welshfella):

just isolate x on its own by adding 14 to each side of the equation and you have it

OpenStudy (logan13):

a

OpenStudy (pinklion23):

do 5-(-14) thrn you will get x

OpenStudy (logan13):

a

OpenStudy (welshfella):

not a

OpenStudy (welshfella):

what is 14 + 5 ?

OpenStudy (pinklion23):

hold on

OpenStudy (logan13):

a

OpenStudy (pinklion23):

yes do that then what do you get

OpenStudy (pinklion23):

14+5 is your answer

OpenStudy (welshfella):

-14 + 14 + x = 5 + 14 x = 5 + 14

OpenStudy (pinklion23):

@logan13 all you have to do is 14 +5 and then you get your answer it is simple :)

OpenStudy (pinklion23):

so what do you get

OpenStudy (pinklion23):

silly logan i know you are hear it say 2 are viewing you and me

OpenStudy (pinklion23):

logan do you know what the answer is

OpenStudy (logan13):

a

OpenStudy (pinklion23):

now really you have a smart score of 66 i know you can figure out 14+5 you silly goose :) now i cant give out the answer or i might get in trouble okay well i hope you find it out